The base-path of a baseball diamond form a square. If it is 90ft from home to first, how far does the catcher have to throw to catch someone stealing second base?
Answer by checkley71(8403)

THE DIAGONAL OF THIS SQUARE IS:
90^2+90^2=D^2
8100+8100=D^2
16,200=D^2
D=SQRT16,200
D=127.28 FEET FROM HOME PLATE TO SECOND BASE.